### v2

@@ Acknowledgments
Jean-Marie Lemercier (jeanmarie.lemercier@uni-hamburg.de) received an M.Eng in Electrical Engineering in 2019 from Ecole Polytechnique, Paris, France. In 2020, he received a M.Sc. in Communications and Signal Processing from Imperial College London, London, United Kingdom. He is currently a PhD student in the Signal Processing group at Universität Hamburg under the supervision of Prof. Dr.-Ing. Timo Gerkmann. His research interests span machine learning-based speech enhancement and dereverberation for hearing devices applications. Recent works also include the design and analysis of diffusion-based generative models for various speech restoration tasks. He is a Student Member of IEEE.
@@
Julius Richter (julius.richter@uni-hamburg.de) received a B.Sc. and M.Sc. in Electrical Engineering in 2017 and 2019 from the Technical University of Berlin, Germany. He is currently a PhD student in the Signal Processing group at Universität Hamburg under the supervision of Prof. Dr.-Ing. Timo Gerkmann. His research interests include deep generative models and multi-modal learning with applications to audio-visual speech processing. He is a Student Member of IEEE.
@@
Simon Welker (simon.welker@uni-hamburg.de) received a B.Sc. in Computing in Science (2019) and an M.Sc. in Bioinformatics (2021) from Universität Hamburg, Germany. He is currently a PhD student under the supervision of Prof. Dr.-Ing. Timo Gerkmann (Signal Processing, Universität Hamburg) and Prof. Dr. Dr. Henry Chapman (Center for Free-Electron Laser Science, DESY, Hamburg), researching machine learning techniques for solving inverse problems that arise in speech processing and X-ray imaging.
@@
Eloi Moliner (eloi.moliner@aalto.fi) received his B.Sc. degree in Telecommunications Technologies and Services Engineering from the Polytechnic University of Catalonia, Spain, in 2018 and his M.Sc. degree in Telecommunications Engineering from the same university in 2021. He is currently a doctoral candidate at the Acoustics Lab of Aalto University in Espoo, Finland. His research interests include digital audio restoration and audio applications of machine learning. He is the winner of the Best Student Paper Award of the 2023 IEEE ICASSP conference.
@@
Vesa Välimäki (vesa.valimaki@aalto.fi) is a Full Professor of audio signal processing and Vice Dean for Research at Aalto University, Espoo, Finland. He received his D.Sc. degree from the Helsinki University of Technology in 1995. In 1996, he was a Postdoctoral Research Fellow at the University of Westminster, London, UK. In 2008–2009, he was a visiting scholar at Stanford University. He is a Fellow of the IEEE, of the Audio Engineering Society, and of the Asia-Pacific Artificial Intelligence Association. Prof. Välimäki is the Editor-in-Chief of the Journal of the Audio Engineering Society.
@@
Timo Gerkmann (timo.gerkmann@uni-hamburg.de) is a Professor for Signal Processing with the Universität Hamburg, Hamburg, Germany. He has held positions with Technicolor Research & Innovation, University of Oldenburg, Oldenburg, Germany, KTH Royal Institute of Technology, Stockholm, Sweden, Ruhr-Universität Bochum, Bochum, Germany, and Siemens Corporate Research, Princeton, NJ, USA. His research interests include statistical signal processing and machine learning for speech and audio applied to communication devices, hearing instruments, audio-visual media, and human-machine interfaces. He was the recipient of the VDE ITG award 2022. He served in the IEEE Signal Processing Society Technical Committee on Audio and Acoustic Signal Processing and is currently a Senior Area Editor of the IEEE/ACM Transactions on Audio, Speech and Language Processing.

### v3

@@ Acknowledgments
Jean-Marie Lemercier (jeanmarie.lemercier@uni-hamburg.de) received an M.Eng in Electrical Engineering in 2019 from Ecole Polytechnique, Paris, France. In 2020, he received a M.Sc. in Communications and Signal Processing from Imperial College London, London, United Kingdom. He is currently a PhD student in the Signal Processing group at Universität Hamburg, 20148 Hamburg, Germany. For his work on speech enhancement with diffusion-based generative models, he was the recipient of the Association for Electrical, Electronic, and Information Technologies Informationstechnische Gesellschaft (VDE ITG) Award 2024. He received the Best Student Paper Award at IWAENC 2024. His research interests span machine learning-based speech enhancement and dereverberation for hearing devices applications. Recent works also include the design and analysis of diffusion-based generative models for various speech restoration tasks. He is a Student Member of IEEE.
@@
Julius Richter (julius.richter@uni-hamburg.de) received a B.Sc. and M.Sc. in Electrical Engineering in 2017 and 2019 from the Technical University of Berlin, Germany. He is currently a PhD student in the Signal Processing group at Universität Hamburg, 20148 Hamburg, Germany. For his work on speech enhancement with diffusion-based generative models, he was the recipient of the Association for Electrical, Electronic, and Information Technologies Informationstechnische Gesellschaft (VDE ITG) Award 2024. His research interests include deep generative models and multi-modal learning with applications to audio-visual speech processing. He is a Student Member of IEEE.
@@
Simon Welker (simon.welker@uni-hamburg.de) received a B.Sc. in Computing in Science (2019) and an M.Sc. in Bioinformatics (2021) from Universität Hamburg, Germany. He is currently a Ph.D. student in the Signal Processing Group, Universität Hamburg, 20148 Hamburg, Germany, and the Coherent Imaging Division, Center for Free-Electron Laser Science, Deutsches Elektronen-Synchrotron DESY, 22607 Hamburg, Germany. For his work on speech enhancement with diffusion-based generative models, he was the recipient of the Association for Electrical, Electronic, and Information Technologies Informationstechnische Gesellschaft (VDE ITG) Award 2024. His research interests include machine learning techniques for solving inverse problems that arise in speech processing and X-ray imaging. He is a Student Member of IEEE.
@@
Eloi Moliner (eloi.moliner@aalto.fi) received his B.Sc. degree in telecommunications technologies and services engineering in 2018 and his M.Sc. degree in telecommunications engineering in 2021 from the Polytechnic University of Catalonia, Spain. He is currently a Ph.D. candidate at the Acoustics Lab of Aalto University, 02150 Espoo, Finland. He is the winner of the Best Student Paper Award of the 2023 IEEE ICASSP conference. He received the Best Student Paper Award at IWAENC 2024. His research interests include digital audio restoration and audio applications of machine learning.
@@
Vesa Välimäki (vesa.valimaki@aalto.fi) received his D.Sc. degree in electrical engineering from the Helsinki University of Technology in 1995. In 1996, he was a post-doctoral research fellow at the University of Westminster, London, U.K. In 2008–2009 he was a visiting scholar at Stanford University. He currently is a full professor of audio signal processing and vice dean for Research at Aalto University, 02150 Espoo, Finland. He is a fellow of the Audio Engineering Society and of the Asia-Pacific Artificial Intelligence Association, and is the editor-in-chief of the Journal of the Audio Engineering Society. His research interests include the application of machine learning and signal processing to audio technology. He is a Fellow of IEEE.
@@
Timo Gerkmann (timo.gerkmann@uni-hamburg.de) received his Dr.-Ing. degree in electrical engineering and information sciences from Ruhr-Universität Bochum, Germany. He is a professor for signal processing with the Universität Hamburg, 20148 Hamburg, Germany. He was the recipient of the Association for Electrical, Electronic, and Information Technologies Informationstechnische Gesellschaft (VDE ITG) Award 2022. He served in the IEEE Signal Processing Society Technical Committee on Audio and Acoustic Signal Processing and is currently a senior area editor of the IEEE/ACM Transactions on Audio, Speech and Language Processing. His research interests include statistical signal processing and machine learning for speech and audio applied to communication devices, hearing instruments, audio-visual media, and human-machine interfaces. He is a Senior Member of IEEE.